OpenStudy (anonymous):

if alpha=(x1;y1) and beta=(x2;y2) then (x1;y1)+(x2;y2)=(x1+x2+1;y1+y2-2) and a*(x1;y1)=(ax1+a-1;y1+y2-2). prove that a(alpha+beta)=a(alpha)+a(beta)

OpenStudy (anonymous):

this are vector

OpenStudy (turingtest):

according to the rule given:\[a\langle x_1,y_1\rangle=\langle ax_1+a-1,y_1+y_2-2\rangle\]is that right?

OpenStudy (turingtest):

seems like there might be a typo...please double-check

OpenStudy (anonymous):

no is like this .

OpenStudy (turingtest):

so we know that\[\alpha+\beta=\langle x_1,y_1\rangle=\langle x_2,y_2\rangle=\langle x_1+x_2+1,y_1+y_2-2\rangle=\langle z_1,z_2\rangle\]and\[a\alpha=a\langle x_1,y_1\rangle=\langle ax_1+a-1,y_1+y_2-2\rangle\]\[a\beta=a\langle x_2,y_2\rangle=\langle ax_2+a-1,y_1+y_2-2\rangle\]hm... this is strange how beta has y1 in it, which makes it depend on alpha are you *sure* there are no typos?

OpenStudy (anonymous):

it is just like this no typus